### v1

# UniCon: A Unified System for Efficient Robot Learning Transfers

Abstract
Deploying learning-based controllers across heterogeneous robots is challenging due to platform differences, inconsistent interfaces, and inefficient middleware. To address these issues, we present UniCon, a lightweight framework that standardizes states, control flow, and instrumentation across platforms. It decomposes workflows into execution graphs with reusable components, separating system states from control logic to enable plug-and-play deployment across various robot morphologies. Unlike traditional middleware, it prioritizes efficiency through batched, vectorized data flow, minimizing communication overhead and improving inference latency. This modular, data-oriented approach enables seamless sim-to-real transfer with minimal re-engineering. We demonstrate that UniCon reduces code redundancy when transferring workflows and achieves higher inference efficiency compared to ROS-based systems. Deployed on over 12 robot models from 7 manufacturers, it has been successfully integrated into ongoing research projects, proving its effectiveness in real-world scenarios.
Figure 1: Representative use cases of UniCon: Left: synchronized and reusable locomotion across heterogeneous robots. Middle: Modular interoperation of RL policies with VR teleoperation. Right: Real‑to‑sim data recording and analysis for diagnosing transfer gaps. Data and control flow are standardized across platforms, reducing integration effort and improving efficiency.
1 Introduction
Deploying learning‑based controllers across heterogeneous robots remains costly due to platform differences and tightly coupled interfaces. These incompatibilities force developers to repeatedly re-implement the same functionalities for each combination of algorithms, robots and simulators. While domain randomization and related sim-to-real techniques (randomization1; randomization2) shrink the dynamics gap, they do not simplify this last-mile integration. Widely used middleware such as ROS and ROS 2 (ros; ros2) attempt to provide a standardized communication layer between customized components, but their general‑purpose design introduces latency and overhead that hinder high‑frequency policy inference (pytorch; Ansel2024PyTorch2F; onnx). At the same time, robot manufacturers complicate the ecosystem further by shipping inconsistent interfaces, making integration and transferability difficult.
To address above challenges, we present UniCon, a lightweight framework for robot learning deployments that provides the unifying infrastructure needed to standardize states, control flow, and instrumentation across platforms. UniCon decomposes workflows into execution graphs with reusable basic units, and separates system states – such as proprioceptive, exteroceptive, and control signals – from the logic that consumes them. This enables plug‑and‑play deployment across diverse morphologies including quadrupeds, humanoids, and manipulators, while unifying hardware with simulators for seamless transfer.
Compared to ROS, our architecture adopt a data-oriented modular paradigm with separation between runtime states and control logic. To maximize efficiency, we favor single-threaded computing (ros2compose) over multiprocessing, and vectorized data flow over message passing. On the other hand, interoperability with existing software is still guaranteed by extending the underlying states storage to multiple communication backends (zeromq2010; ddsfoundation2025). We show that this reduces code redundancy and improves inference latency in our cross-embodiment experiments.
Our contributions are threefold:
  • Unified control framework: We introduce UniCon, a lightweight and efficient framework that bridges mainstream simulators and physical hardware for robot learning.
  • Modular, data-oriented design: We demonstrate that UniCon enables efficient sim-to-real transfer across diverse robot embodiments with little to no re-engineering effort.
  • Improved inference efficiency: We show that UniCon achieves higher inference efficiency compared to ROS-based and ad-hoc framework stacks by reducing interfacing and communication overhead.
UniCon has already been deployed on over 12 robot models spanning 7 manufacturers, and has supported 3 research projects ongoing in the Shanghai Artificial Intelligence Laboratory, demonstrating its versatility in real research scenarios. We release our framework as open-source at https://github.com/creeperlin/unicon.
2 Framework
A key obstacle in sim‑to‑real and cross-embodiment transfer is the tight coupling of system data flows with customized control logic. Middleware such as ROS enforce a message passing pattern but incurs latency, making it unsuitable for learning‑based controllers. Our design instead adopts a modular, data‑oriented paradigm that cleanly separates states from logic, emphasizing reusability and efficiency. This approach contrasts with object‑oriented programming practices (10.5555/1407387), where data and routines are encapsulated together, and instead aligns more closely with the Entity Component System architecture widely used in the game industry (overwatch2017gdc) and modern simulation engines (shacklett23madrona).
Figure 2: Architecture of UniCon: (a) global system states with switchable storage backends; (b) modular control blocks covering platform and inference; (c) control flow graph primitives for workflow composition; and (d) unified integration with simulators and hardware.
2.1 Vectorized States
We formulate all runtime states of a robot workflow as global array objects of numerical data types under different labels. This comes from the fact that the data passed in structured messages often take a fixed format for each channel if the hardware participants remains constant. Therefore, rather than transmitting full self‑describing objects, only the essential multidimensional data are exchanged, eliminating unnecessary encapsulation and aligning with contiguous memory layouts favored by simulators and hardware interfaces.
Typically, the states includes robot sensory inputs such as motor position and velocities, as well as orientation and angular velocity reported by the IMU. Similarly, motor commands such as position and torque controls are also states which are written to by controllers and read from by hardware interfaces. In the case where state definitions of custom components differ from the actual hardware, transparent mappings with efficient vectorized indexing operations is inserted to align properties such as joint order.
2.2 Modularized Logic
Following the vectorized formulation, we break down workflows into reusable Control Blocks (CB) bridged by global states. For example, operations like receiving hardware states, executing policy inference, handling user inputs, and sending hardware commands are split into individual blocks and connected into an execution graph through standardized control flow. This grants users the ability to shift between controllers, policies and hardware by switching between CBs, without changing the rest of the graph.
Formally, we assume the global state space $S=\{s_{a},s_{b},\dots,s_{n}\},s_{i}\in{\mathcal{R}}^{k_{i}}$ is the common domain of all operating logic. Then we define a CB as a pure function with regard to $S$. It can read a subset of $S$ as inputs and outputs updated values of some other elements in $S$, and a boolean $r$ indicating its termination.
$f:S_{in}\mapsto S_{out}\times\{0,1\},S_{in},S_{out}\subseteq S.$ (1)
Stepping the function results in a sequence of calls:
$T_{f}=(f_{1},f_{2},\dots,f_{n}).$ (2)
A workflow is then constructed by nesting the CBs, forming an execution graph. We provide basic control flow components in similar fasion to functional programming (Hughes1989WhyFP) for composing customized workflows. For example, loop terminates the inner block when the predicate is satisfied, zip executes multiple blocks in sequence, and chain exhausts nested blocks in order until they all halts.
$\operatorname{loop}(T_{g},p)$ $=(g_{1},g_{2},\dots,g_{k}),p(k)=1,$ (3)
$\operatorname{zip}(T_{g},T_{h})$ $=((g_{1},h_{1}),(g_{2},h_{2}),\dots,(g_{n},h_{n})),$
$\operatorname{chain}(T_{g},T_{h})$ $=(g_{1},\dots,g_{m},h_{1},\dots,h_{n}).$
This enables configurations via text formats and possibly GUIs, while also leaves directives for code generation.
2.3 Software Interoperability
Communications.  UniCon connects its state storage backend to external clients with consistent semantics: writing to arrays publishes data, reading retrieves the latest values. Multiple backends provide flexible performance: lock‑free local or shared memory for zero‑copy access (Poehnl2019Iceoryx), message queues for distributed setups (zeromq2010), and ROS bridges for legacy integration.
Platform support.  Simulation and hardware platforms are integrated uniformly as state readers and writers. Each platform adapter provides three callable blocks: recv for updating states, send for sending controls, and close for teardown, while automatically aligning parameters such as PID gains and DoF limits to ensure consistent control semantics (Chitta2017roscontrolAG). Supported platforms are summarized in Table 1.
Peripherals.  We also incoporate frequently used peripheral tools and devices as plug-and-play CBs. Input devices such as joysticks (linuxinput2025) and VR (vuer) are supported with multiple implementations for coverage, while sensors including RGBD (realsense) cameras and LiDARs are connected either through platform proxy or standalone clients. Visualizers such as Foxglove and Meshcat (qin2023anyteleop) are also ready to be wired to selected states.
2.4 Programming Interface
UniCon exposes states as typed NumPy (harris2020array) arrays, enabling users to perform vectorized operations in Python. Workflows can be defined either programmatically or textually, with user logic seamlessly integrated via parameter-less boolean functions following our CB design. This causes minimal intrusion while allowing users to fully customize their control flow as needed.
2.5 Workflow Transfers
UniCon enables seamless tranfer between three workflows: Sim-to-Sim validation, Sim-to-Real deployment, and Real-to-Sim instrumentation, all operating on a unified data abstraction. Sim-to-Sim and Sim-to-Real are trivial since the system CBs remain interchangeable within the execution graph. For real deployments, additional safety blocks can be inserted to handle startup, shutdown, and runtime monitoring without altering the rest of the graph.
Real-to-Sim uses recorder and replay blocks on physical states and built‑in metrics to quantify the reality gap. For open‑loop and repeatable rollouts, mean squared error (MSE) performs step‑wise comparison since inputs are exactly replayed. For closed‑loop controls like locomotion, we propose to unfold the MSE loss to prevent misalignments:
$\mathcal{L}=\frac{1}{n}\min_{j}\sum_{i=0}^{n}\|T_{i+j}-\hat{T}_{i}\|_{2}+\frac{1}{n}\min_{j}\sum_{i=0}^{n}\|T_{i}-\hat{T}_{i+j}\|_{2}.$ (4)
3 Evaluation
Table 1: Supported simulators and hardware platforms in UniCon.
Simulators Humanoids Quadrupeds
IsaacGym (isaacgym) Unitree H1 (unitree) Unitree A1
IsaacSim (nvidia2023isaacsim) Unitree G1 Unitree Aliengo
IsaacLab (orbit) Unitree H1-2 Unitree Go2
MuJoCo (mujoco) Fourier GR1 (fftai)
MuJoCo MJX (mujoco_playground_2025) Fourier N1 Manipulators
PyBullet (pybullet) AgiBot X2 (agibot2023x2) ARX (arxroboticsx2025github)
Webots (Michel2004CyberboticsLW) Dobot Atom (atom) ROHand (rohand2025oymotion)
Genesis (Genesis) PND Adam (pnd) Unitree Dex3-1
Gazebo (Koenig2004DesignAU) AzureLoong (oghr2)
Newton (newton2025github) Booster T1 (t1)
We evaluate our framework on deployment tasks across quadrupeds, humanoids, and manipulators, focusing on transfer effort, inference efficiency and workflow usability.
3.1 Transfer Effort
We target the locomotion task for humanoids and quadrupeds where policies with the same training paradigm are deployed on multiple platforms (xue2025unified). Without a framework, the inference often need to be re‑implemented due to vendor‑specific stacks. We report the Source Lines of Code (SLOC) required to transfer the original workflow on Unitree H1 to the MuJoCo simulator and other robots. As shown in Table 2, the Ad‑hoc code needs substantial efforts to transfer, especially for the PND Adam model with its completely different APIs, while UniCon remains lightweight and requires no extra effort when switching deployment targets.
Table 2: Comparison of transfer effort across workflows (code changes, in SLOC).
Framework Components Inference workflows
H1 Sim G1 Adam NN H1 Sim G1 Adam
Ad-hoc 852 (original code base for H1) 0 507 627 770
UniCon (ours) 393 382 35 394 350 0 0 0 0
3.2 Framework Efficiency
We measure the framework efficiency in the form of operation latency overhead. The evaluation runs on the same H1 model with identity joint position control ($q_{d}=q$) at 50Hz in place of actual inference. For this robot, system states are refreshed at around 500Hz by the SDK (unitreeSDK2python2025). The synchronous (Sync) setup queries the global buffer directly at each control cycle, while the asynchronous (Async) setup uses a callback to receive the latest states. In the table, Recv denotes the time to obtain the latest system states, Send denotes the time to transmit the resulting control signal, and End-to-end measures the latency between sending a control signal and the timestamp of the dependent states. Because the state timestamp originates from a monotonic clock on another onboard system, we manually align clocks with an inferred time origin for this row.
Results in Table 3 show that UniCon introduces near‑zero overhead compared to SDK implementations and significantly outperforms ROS. This efficiency arises from binding global state buffers into the data reader/writer and compiling (pybind11github; RoboJuDo) alongside the SDK to eliminate redundant data copies. During actual inference, control blocks can be arranged so that the critical path contains only the data producers required by the policy, deferring non‑critical communication and synchronization.
[…]
Operation Frameworks
SDK (Sync) SDK (Async) ROS 2 UniCon (ours)
Recv 1261 $\pm$ 150 - 50 $\pm$ 88 63 $\pm$ 15
Send 1011 $\pm$ 72 1001 $\pm$ 65 587 $\pm$ 179 190 $\pm$ 45
End-to-end11footnotemark: 1 1805 $\pm$ 1559 1447 $\pm$ 936 1658 $\pm$ 882 732 $\pm$ 749
11footnotemark: 1 Time between send and state timestamp, applying an inferred offset due to relative system clocks.
3.3 Workflow Usability
While ROS is the go-to choice for building complex robotic application, UniCon offers comparable usability with built-in connection to mainstream sensors, input devices, motion libraries and messaging layers including ROS itself. We demonstrate this by extending the inference pipeline to wholebody control, incorporating inverse kinematics (Buss2004IntroductionTI; pink), dexterous hand retargeting (qin2023anyteleop), and VR teleoperation (Cheng2024OpenTeleVisionTW), all modularized and distributed across onboard and LAN workstations.
We further showcase real‑to‑sim analysis by comparing recorded and replayed trajectories across systems. Using the built‑in analyzer, discrepancies can be measured element‑wise and frame‑wise to identify sources of the reality gap. As shown in Figure 3, a bipedal policy stable in the IsaacGym simulation but failing on the A1 robot exhibits significant deviation in the rear left calf joint, leading to overheating and eventual loss of balance. In contrast, the Go2 model shows a smaller gap under similar settings.
Refer to caption Refer to caption Refer to caption
Figure 3: Real-to-sim analysis of inference trajectories. Left & Middle: Reality gap quantified per joint position using built-in metrics, showing deviations in the A1’s rear left calf joint. Right: Stable Go2 standing (top) and A1 just before fallover (bottom).
4 Related Work
4.1 Learning-based Robot Controllers
Learning-based controllers, particularly reinforcement learning (RL), have enabled locomotion across quadrupeds (lee2020learning; smith2022; rma; hoeller2024anymal), wheeled robots (lee2024learning), and humanoids (gu2024advancing; xue2025unified; wang2025beamdojo; huang2025learning), achieving agile gaits (amp-wu; he2024agile), terrain adaptation (tert; parkour1; parkour2; Lai2024WorldMP), whole-body control (cheng2024express; ji2024exbody2), and motion tracking (amp; yin2025unitrackerlearninguniversalwholebody). These advances highlight the promise of RL for scalable locomotion, but also the challenge of sim-to-real transfer: policies trained in simulation often fail to generalize due to discrepancies in dynamics and perception (comparesim; da2025survey).
To address this, domain randomization (randomization1; OpenAI2019SolvingRC; randomization2; campanaro2022) and adaptation (minimizing) improve robustness, while real-to-sim system identification (SysId) methods calibrate simulators with real-world data (dong2024easi; he2025asap; xu2025deal). UniCon complements these strategies by introducing a normalization pipeline and hardware-agnostic interface abstraction, enabling cross-platform inference, data gathering and analysis.
4.2 Robot Application Middleware
Robot application middleware provides critical abstractions and communication infrastructure for integrating perception, planning, and control components across heterogeneous robotic systems. Frameworks such as the Robot Operating System (ROS) have become de facto standards in academia and industry due to their modularity, device driver ecosystem, and support for distributed computing (ros; ros2; Chitta2012MoveItT). Similarly, PX4 (Meier2015PX4AN) offers a full-stack flight control system for aerial robotics, facilitating high-performance real-time control with rich simulation integration.
While these middleware platforms simplify software development and system orchestration, they generally lack standardized abstractions for cross-platform strategy deployment or sim-to-real policy learning. UniCon complements these frameworks by introducing a domain-aware control layer specifically tailored for learning-based policy modularization and transferability. Rather than replacing existing middleware, UniCon operates orthogonally – interfacing with existing environments to unify and reuse control logic across embodiments.
5 Conclusion
UniCon offers a unified control layer in robot learning deployments by separating vectorized global states from modular control blocks, greatly reducing hardware transfer effort and enabling low-latency inference. Its successful use across robots and research projects show strong potentials for broader community adoption. Future works include multi‑language APIs and code generation to broaden platform coverage.
Acknowledgments
The Shanghai Jiao Tong University team is partially supported by National Natural Science Foundation of China (62322603). We also thank Fourier Intelligence and ByteDance Seed for their early support with hardware and experiments.

### v2

@@
 At the same time, robot manufacturers complicate the ecosystem further by shipping inconsistent interfaces, making integration and transferability difficult.
-To address above challenges, we present UniCon, a lightweight framework for robot learning deployments that provides the unifying infrastructure needed to standardize states, control flow, and instrumentation across platforms.
+To address these challenges, we present UniCon, a lightweight framework for robot learning deployments that provides the unifying infrastructure needed to standardize states, control flow, and instrumentation across platforms.
 UniCon decomposes workflows into execution graphs with reusable basic units, and separates system states – such as proprioceptive, exteroceptive, and control signals – from the logic that consumes them.
 This enables plug‑and‑play deployment across diverse morphologies including quadrupeds, humanoids, and manipulators, while unifying hardware with simulators for seamless transfer.
-Compared to ROS, our architecture adopt a data-oriented modular paradigm with separation between runtime states and control logic.
+Compared to ROS, our architecture adopts a data-oriented modular paradigm with separation between runtime states and control logic.
 To maximize efficiency, we favor single-threaded computing (ros2compose) over multiprocessing, and vectorized data flow over message passing.
@@
   • Improved inference efficiency: We show that UniCon achieves higher inference efficiency compared to ROS-based and ad-hoc framework stacks by reducing interfacing and communication overhead.
-UniCon has already been deployed on over 12 robot models spanning 7 manufacturers, and has supported 3 research projects ongoing in the Shanghai Artificial Intelligence Laboratory, demonstrating its versatility in real research scenarios.
+UniCon has already been deployed on over 12 robot models spanning 7 manufacturers, and has supported 3 research projects in the Shanghai Artificial Intelligence Laboratory, demonstrating its versatility in real research scenarios.
 We release our framework as open-source at https://github.com/creeperlin/unicon.
@@
 A key obstacle in sim‑to‑real and cross-embodiment transfer is the tight coupling of system data flows with customized control logic.
-Middleware such as ROS enforce a message passing pattern but incurs latency, making it unsuitable for learning‑based controllers.
+Middleware such as ROS enforces a message passing pattern but incurs latency, making it unsuitable for learning‑based controllers.
 Our design instead adopts a modular, data‑oriented paradigm that cleanly separates states from logic, emphasizing reusability and efficiency.
@@
 We formulate all runtime states of a robot workflow as global array objects of numerical data types under different labels.
-This comes from the fact that the data passed in structured messages often take a fixed format for each channel if the hardware participants remains constant.
+This comes from the fact that the data passed in structured messages often takes a fixed format for each channel if the hardware participants remains constant.
 Therefore, rather than transmitting full self‑describing objects, only the essential multidimensional data are exchanged, eliminating unnecessary encapsulation and aligning with contiguous memory layouts favored by simulators and hardware interfaces.
-Typically, the states includes robot sensory inputs such as motor position and velocities, as well as orientation and angular velocity reported by the IMU.
+Typically, the states include robot sensory inputs such as motor position and velocities, as well as orientation and angular velocity reported by the IMU.
 Similarly, motor commands such as position and torque controls are also states which are written to by controllers and read from by hardware interfaces.
-In the case where state definitions of custom components differ from the actual hardware, transparent mappings with efficient vectorized indexing operations is inserted to align properties such as joint order.
+In the case where state definitions of custom components differ from the actual hardware, transparent mappings with efficient vectorized indexing operations are inserted to align properties such as joint order.
 2.2 Modularized Logic
@@
 A workflow is then constructed by nesting the CBs, forming an execution graph.
-We provide basic control flow components in similar fasion to functional programming (Hughes1989WhyFP) for composing customized workflows.
-For example, loop terminates the inner block when the predicate is satisfied, zip executes multiple blocks in sequence, and chain exhausts nested blocks in order until they all halts.
+We provide basic control flow components in similar fashion to functional programming (Hughes1989WhyFP) for composing customized workflows.
+For example, loop terminates the inner block when the predicate is satisfied, zip executes multiple blocks in sequence, and chain exhausts nested blocks in order until they all halt.
 $\operatorname{loop}(T_{g},p)$ $=(g_{1},g_{2},\dots,g_{k}),p(k)=1,$ (3)
@@
 $\operatorname{chain}(T_{g},T_{h})$ $=(g_{1},\dots,g_{m},h_{1},\dots,h_{n}).$
-This enables configurations via text formats and possibly GUIs, while also leaves directives for code generation.
+This enables configurations via text formats and possibly GUIs, while also leaving directives for code generation.
 2.3 Software Interoperability
@@
 Peripherals.
-We also incoporate frequently used peripheral tools and devices as plug-and-play CBs.
-Input devices such as joysticks (linuxinput2025) and VR (vuer) are supported with multiple implementations for coverage, while sensors including RGBD (realsense) cameras and LiDARs are connected either through platform proxy or standalone clients.
+We also incorporate frequently used peripheral tools and devices as plug-and-play CBs.
+Input devices such as joysticks (linuxinput2025) and VR (vuer) are supported with multiple implementations for coverage, while sensors including RGBD (realsense) cameras and LiDARs are connected either through platform proxy or standalone clients; high-bandwidth image data are handled through their native interfaces and referenced via lightweight handles to preserve efficiency.
 Visualizers such as Foxglove and Meshcat (qin2023anyteleop) are also ready to be wired to selected states.
@@
 2.5 Workflow Transfers
-UniCon enables seamless tranfer between three workflows: Sim-to-Sim validation, Sim-to-Real deployment, and Real-to-Sim instrumentation, all operating on a unified data abstraction.
+UniCon enables seamless transfer between three workflows: Sim-to-Sim validation, Sim-to-Real deployment, and Real-to-Sim instrumentation, all operating on a unified data abstraction.
 Sim-to-Sim and Sim-to-Real are trivial since the system CBs remain interchangeable within the execution graph.
@@
 For open‑loop and repeatable rollouts, mean squared error (MSE) performs step‑wise comparison since inputs are exactly replayed.
-For closed‑loop controls like locomotion, we propose to unfold the MSE loss to prevent misalignments:
-$\mathcal{L}=\frac{1}{n}\min_{j}\sum_{i=0}^{n}\|T_{i+j}-\hat{T}_{i}\|_{2}+\frac{1}{n}\min_{j}\sum_{i=0}^{n}\|T_{i}-\hat{T}_{i+j}\|_{2}.$ (4)
+For closed‑loop controls like locomotion, we propose to unfold the MSE loss to prevent temporal misalignments:
+$\mathcal{L}=\frac{1}{n}\min_{j}\sum_{i=0}^{n}\|T_{i+j}-\hat{T}_{i}\|_{2}+\frac{1}{n}\min_{j}\sum_{i=0}^{n}\|T_{i}-\hat{T}_{i+j}\|_{2},$ (4)
+where $T$ represents the real robot trajectory, $\hat{T}$ represents the simulated trajectory, and $j$ is the temporal shift minimized to find the best alignment between the two sequences.
 3 Evaluation
@@
 Genesis (Genesis) PND Adam (pnd) Unitree Dex3-1
-Gazebo (Koenig2004DesignAU) AzureLoong (oghr2)
-Newton (newton2025github) Booster T1 (t1)
+Gazebo (Koenig2004DesignAU) AzureLoong (oghr2) Inspire RH56
+Newton (newton2025github) Booster T1 (t1) XHand1
 We evaluate our framework on deployment tasks across quadrupeds, humanoids, and manipulators, focusing on transfer effort, inference efficiency and workflow usability.
@@
 We target the locomotion task for humanoids and quadrupeds where policies with the same training paradigm are deployed on multiple platforms (xue2025unified).
-Without a framework, the inference often need to be re‑implemented due to vendor‑specific stacks.
+Without a framework, the inference often needs to be re‑implemented due to vendor‑specific stacks.
 We report the Source Lines of Code (SLOC) required to transfer the original workflow on Unitree H1 to the MuJoCo simulator and other robots.
@@
 Because the state timestamp originates from a monotonic clock on another onboard system, we manually align clocks with an inferred time origin for this row.
-Results in Table 3 show that UniCon introduces near‑zero overhead compared to SDK implementations and significantly outperforms ROS.
+Results in Table 3 show that UniCon introduces near‑zero overhead compared to Python SDK implementations and significantly outperforms ROS.
 This efficiency arises from binding global state buffers into the data reader/writer and compiling (pybind11github; RoboJuDo) alongside the SDK to eliminate redundant data copies.
@@
 3.3 Workflow Usability
-While ROS is the go-to choice for building complex robotic application, UniCon offers comparable usability with built-in connection to mainstream sensors, input devices, motion libraries and messaging layers including ROS itself.
+While ROS is the go-to choice for building complex robotic applications, UniCon offers comparable usability with built-in connection to mainstream sensors, input devices, motion libraries and messaging layers including ROS itself.
 We demonstrate this by extending the inference pipeline to wholebody control, incorporating inverse kinematics (Buss2004IntroductionTI; pink), dexterous hand retargeting (qin2023anyteleop), and VR teleoperation (Cheng2024OpenTeleVisionTW), all modularized and distributed across onboard and LAN workstations.
@@
 4.2 Robot Application Middleware
-Robot application middleware provides critical abstractions and communication infrastructure for integrating perception, planning, and control components across heterogeneous robotic systems.
+Robot application middleware provides critical abstractions and communication infrastructure for integrating perception, planning, and control components across heterogeneous robotic systems (Gerkey2003ThePP; Metta2006YARPYA; Bruyninckx2001OpenRC; Utz2002MiroM).
 Frameworks such as the Robot Operating System (ROS) have become de facto standards in academia and industry due to their modularity, device driver ecosystem, and support for distributed computing (ros; ros2; Chitta2012MoveItT).
@@
 UniCon offers a unified control layer in robot learning deployments by separating vectorized global states from modular control blocks, greatly reducing hardware transfer effort and enabling low-latency inference.
-Its successful use across robots and research projects show strong potentials for broader community adoption.
-Future works include multi‑language APIs and code generation to broaden platform coverage.
+Its successful use across robots and research projects shows strong potential for broader community adoption.
+Future work includes multi‑language APIs and code generation to broaden platform coverage.
 Acknowledgments
-The Shanghai Jiao Tong University team is partially supported by National Natural Science Foundation of China (62322603).
+The SJTU team is supported by National Natural Science Foundation of China (62322603), Shanghai Municipal Science and Technology Major Project (2025SHZDZX025D08) and Shanghai Artificial Intelligence Laboratory.
 We also thank Fourier Intelligence and ByteDance Seed for their early support with hardware and experiments.